# SpecCLIP: Aligning and Translating Spectroscopic Measurements for Stars

### v1

@@ 1 Introduction
  • We demonstrate that our unified model enables both in-modality and cross-modality search, and cross-modality prediction.
@@ 2.3 Contrastive Learning with Decoders
  • Two in-modality decoders reconstruct each spectrum from its projected embedding;
  • Two cross-modality decoders predict one modality of spectrum from the other.
@@
  • CLIP-p: Adds only cross-modality prediction decoders to the baseline.
@@ 2.4 Loss Functions during Contrastive Training
Together, these losses ensure that the model aligns LRS and XP modalities in a shared space (via contrastive loss), retains informative modality-specific content (via reconstruction), and supports spectrum translation across modalities (via prediction). In CLIP-split, this is achieved through a disentangled embedding structure with explicit shared and non-shared components, while CLIP-pr uses a unified embedding vector for all tasks.
@@ 3.2 Sample Selection for Parameter Estimation
Table 1: Comparison of Model Performance (standard deviation of the residuals $\sigma$ and coefficient of determination $R^{2}$) for different models evaluating on the held-out test datasets.
@@
$\mathrm{[\alpha/Fe]}$ 0.023 / --2.376 0.022 / 0.901 0.021 / 0.906 0.020/0.910 0.020 / 0.909 0.020 / 0.909 0.020 / 0.910
$\mathrm{[C/Fe]}$ 0.039 / 0.000 0.039 / 0.784 0.038 / 0.806 0.038/0.805 0.038 / 0.805 0.038 / 0.805 0.038 / 0.806
$\mathrm{[N/Fe]}$ 0.054 / --6.346 0.052 / 0.630 0.050 / 0.659 0.050/0.660 0.050 / 0.660 0.050 / 0.660 0.050 / 0.661
$\mathrm{[Al/Fe]}$ 0.048 / --2.404 0.049 / 0.703 0.046 / 0.742 0.046/0.742 0.046 / 0.741 0.046 / 0.742 0.047 / 0.736
$\mathrm{[Ca/Fe]}$ 0.031 / --8.202 0.031 / 0.691 0.029 / 0.713 0.029/0.716 0.029 / 0.714 0.029 / 0.716 0.030 / 0.708
$\mathrm{[Mg/Fe]}$ 0.031 / --4.301 0.032 / 0.867 0.031 / 0.878 0.031/0.877 0.031 / 0.878 0.031 / 0.878 0.031 / 0.879
$\mathrm{[Si/Fe]}$ 0.030 / --5.773 0.029 / 0.798 0.028 / 0.807 0.028/0.808 0.028 / 0.809 0.028 / 0.811 0.028 / 0.807
$\mathrm{[Ti/Fe]}$ 0.058 / --0.339 0.058 / 0.528 0.056 / 0.546 0.056/0.548 0.056 / 0.545 0.056 / 0.547 0.057 / 0.543
$\mathrm{[Mn/Fe]}$ 0.034 / --6.203 0.032 / 0.779 0.031 / 0.795 0.031/0.797 0.031 / 0.795 0.031 / 0.797 0.031 / 0.791
$\mathrm{[Ni/Fe]}$ 0.026 / --0.083 0.026 / 0.450 0.025 / 0.486 0.025/0.486 0.025 / 0.485 0.025 / 0.485 0.026 / 0.480
$\mathrm{[O/Fe]}$ 0.052 / --4.127 0.051 / 0.717 0.049 / 0.726 0.049/0.727 0.049 / 0.725 0.049 / 0.726 0.049 / 0.724
$\mathrm{[Cr/Fe]}$ 0.078 / --0.575 0.076 / 0.222 0.075 / 0.234 0.075/0.232 0.075 / 0.233 0.075 / 0.234 0.075 / 0.231
@@
$\Delta\nu$ 1.398/0.938 1.661/0.936 1.550/0.890 1.499/0.913 1.512/0.855 1.555/0.888 1.562/0.946
$\nu_{\rm max}$ 17.568/0.632 23.826/0.569 21.842/0.431 22.093/0.524 22.470/0.337 22.398/0.595 22.444/0.522
Mass ($M_{\odot}$) 0.094/0.554 0.098/0.557 0.088/0.670 0.088/0.649 0.089/0.661 0.089/0.650 0.092/0.647
Radius ($R_{\odot}$) 0.570/0.886 0.662/0.855 0.702/0.841 0.714/0.846 0.772/0.848 0.645/0.862 0.600/0.893
Age (Gyr) 1.561/0.648 1.587/0.640 1.397/0.721 1.347/0.747 1.359/0.749 1.373/0.745 1.394/0.712
$\Delta\Pi$ 28.950/0.877 25.181/0.904 22.556/0.903 22.889/0.897 21.691/0.912 22.332/0.910 22.868/0.918
@@
$\mathrm{[\alpha/Fe]}$ 0.103 / --0.047 0.056 / 0.737 0.049 / 0.774 0.048/0.777 0.049 / 0.773 0.049 / 0.770 0.050 / 0.766
$\mathrm{[C/Fe]}$ 0.194 / 0.073 0.127 / 0.527 0.118 / 0.547 0.118/0.553 0.118 / 0.550 0.117 / 0.549 0.118 / 0.552
$\mathrm{[N/Fe]}$ 0.115 / --4.040 0.077 / 0.643 0.072 / 0.673 0.072/0.676 0.073 / 0.672 0.072 / 0.674 0.073 / 0.671
@@
$E(BP-RP)$ 0.077 / 0.725 0.036 / 0.921 0.036 / 0.926 0.035/0.927 0.036 / 0.925 0.035 / 0.926 0.034 / 0.929
Number of wins (best $\sigma$ or $R^{2}$) 0 0 19 29 15 26 22
@@ 4.1 Model Comparison
In particular, if checked more carefully, models with in-modality reconstruction decoders (e.g., CLIP-r and CLIP-pr compared with CLIP, and CLIP-pr compared with CLIP-p) generally show improved performance, as revealed by the “number of wins” in the table (last row), fairly accounting for the MLP-based downstream models only. This indicates enhanced informativeness of the learned representations for downstream parameter estimation. Overall, these results suggest that the inclusion of in-modality reconstruction decoders improves representation quality in most downstream applications.
We hypothesize that these performance gains stem from the model’s ability to retain both shared and modality-specific (non-shared) information. In the CLIP-pr variant, this is encouraged by jointly optimizing the contrastive loss, cross-modal prediction loss, and in-modality reconstruction loss. This architecture implicitly encourages the embeddings to retain complementary information from each modality.
@@
Interestingly, CLIP-split achieves the highest similarity score overall, even surpassing the baseline CLIP model. It also delivers the best performance on XP $\rightarrow$ LRS spectrum prediction, demonstrating the model’s robustness across modalities.
@@
Figure 2: Comparison between the LAMOST catalog and SpecCLIP models (including the pre-trained LRS model and the LRS branch of the CLIP-split model). From top to bottom: The radial velocity (RV) comparison as a function of the GALAH labels; [Fe/H] comparison as a function of the DESI labels; [Fe/H] comparison as a function of the GALAH labels; and [Fe/H] comparison as a function of the GALAH labels with input spectra shifted to the rest frame using the predicted RVs from the corresponding models in the top row.
For RV, which is inferred using the SBI downstream model, the pre-trained LRS model and CLIP-split model have slightly larger scatter but smaller bias, compared with the LAMOST catalog; For [Fe/H], inferred using MLP downstream models (as with all other figures), the CLIP-split model gives either smaller scatter over the metal-poor region (referring to DESI labels) or overall smaller scatter and bias (referring to GALAH labels). The RV-corrected spectra result in similar [Fe/H] prediction performance, suggesting that the trained MLP models are relatively robust to modest Doppler shifts in the LAMOST LRS spectra. The dashed lines are the one-to-one lines. The numbers is the upper left of each panel are the mean offsets and standard deviation of the residuals.
@@ 4.2.2 Gaia XP
Figure 5: Two examples of in-modality retrieval, cross-modality retrieval, cross-modality prediction, and the LAMOST (Gaia) spectra corresponding to Gaia (LAMOST) in-modality retrieval. The similarity scores are defined in the projected embedding space.
Table 3: Comparison of Model Performance (standard deviation of the residuals $\sigma$ and coefficient of determination $R^{2}$) for pre-trained XP models with different embedding dimensions (256, 343, 512, 768), where 768 is adopted in this paper.
@@ 4.3.1 Spectral Retrieval
Figure 5 shows examples of in-modality and cross-modality retrievals given a specific query spectrum. Retrieval is based on cosine similarity in the (projected) embedding space, using either the full or shared embeddings (for CLIP-split). The search is conducted over a test set of 82,057 spectra, with the query spectrum excluded from the database. In both LAMOST LRS and Gaia XP cases, the retrieved spectra closely resemble the query spectra, indicating that the model has learned well-aligned representations across modalities.
In practice, additional strategies can be used to retrieve spectra using auxiliary catalog links. For example, given a query LAMOST spectrum and a LAMOST-to-Gaia cross-match library, one could first retrieve the top LAMOST matches (in-modality) and then fetch their Gaia counterparts via the library. Alternatively, if the paired Gaia spectrum of the query is known, one could retrieve similar Gaia spectra directly (in-modality), or indirectly by performing a Gaia-to-LAMOST retrieval followed by a database lookup to obtain the corresponding Gaia spectra. While Figure 5 presents only two retrieval use cases, assuming that we know only the information of the query spectrum itself, not its paired other-modal spectrum. The other more elaborate approaches are straightforward extensions.
@@ 4.3.2 Spectral Prediction
SpecCLIP’s cross-modality decoders also support spectral translation, that is, predicting the spectrum in one modality from a spectrum in another. These decoders operate directly on the projected embeddings (shared embeddings in the case of CLIP-split), using learned mappings between modalities.
@@
However, for a subset of sources—particularly in the LRS $\rightarrow$ XP direction—prediction quality deteriorates. Apart from the potential effect of extinction poorly-learned by the trained model, this discrepancy may suggest that the source does not follow the behavior of a typical single star. For instance, it could be an unresolved binary or an otherwise anomalous object. These cases highlight an exciting future direction, using cross-modality prediction error as a basis for anomaly detection.
@@ 5 Discussion
5.1 SBI Performance for [Fe/H]nd Radial Velocity
In our experiments with SBI, we observed that applying NSF to projected embeddings led to degraded performance in the low-iron abundance regime ($\mathrm{[Fe/H]}\lesssim-2$). Similarly, radial velocity predictions from LAMOST LRS exhibited systematic underestimation at large absolute radial velocities; a similar issue broadly discussed in machine learning-based estimators (Ting, 2024).
@@ 5.1 SBI Performance for [Fe/H]nd Radial Velocity
A common assumption is that compression improves downstream performance. However, our Gaia XP foundation model shows a more nuanced behavior. In Table 3, we vary the embedding dimension and find that reducing it from 343 (the original XP spectrum length) to 256 leads to worse performance. In contrast, using embedding dimensions equal to or greater than the input size yields better results.
@@ 5.5 Training Sample Size and Dataset Configuration
Nevertheless, even with this moderate sample size, our models match or outperform state-of-the-art results in the literature (Section 4.2). To maximize data usage, we report metrics based on held-out test sets, while for plots generated using MLP-based downstream models, we combine the training and validation sets for training, with early stopping based on performance on the test set.
@@ 5.7 Fine-tuning Loss Weights
Table 4: Comparison of Model Performance (standard deviation of the residuals $\sigma$ and coefficient of determination $R^{2}$) between masked transformer (MT) and MLP-based ordinary auto-encoder (OAE).
@@
$\mathrm{[\alpha/Fe]}$ 0.023 / --2.376 0.022 / 0.901 0.020 / 0.905
$\mathrm{[C/Fe]}$ 0.039 / 0.000 0.039 / 0.784 0.039 / 0.778
$\mathrm{[N/Fe]}$ 0.054 / --6.346 0.052 / 0.630 0.053 / 0.621
$\mathrm{[Al/Fe]}$ 0.048 / --2.404 0.049 / 0.703 0.049 / 0.693
$\mathrm{[Ca/Fe]}$ 0.031 / --8.202 0.031 / 0.691 0.031 / 0.688
$\mathrm{[Mg/Fe]}$ 0.031 / --4.301 0.032 / 0.867 0.030 / 0.872
$\mathrm{[Si/Fe]}$ 0.030 / --5.773 0.029 / 0.798 0.029 / 0.795
$\mathrm{[Ti/Fe]}$ 0.058 / --0.339 0.058 / 0.528 0.059 / 0.507
$\mathrm{[Mn/Fe]}$ 0.034 / --6.203 0.032 / 0.779 0.033 / 0.761
$\mathrm{[Ni/Fe]}$ 0.026 / --0.083 0.026 / 0.450 0.027 / 0.446
$\mathrm{[O/Fe]}$ 0.052 / --4.127 0.051 / 0.717 0.051 / 0.703
$\mathrm{[Cr/Fe]}$ 0.078 / --0.575 0.076 / 0.222 0.079 / 0.198

### v2

@@ 1 Introduction
  • We demonstrate that our unified model enables both in-modal and cross-modal search, and cross-modal prediction.
@@ 2.3 Contrastive Learning with Decoders
  • Two in-modal decoders reconstruct each spectrum from its projected embedding;
  • Two cross-modal decoders predict one modality of spectrum from the other.
@@
  • CLIP-p: Adds only cross-modal prediction decoders to the baseline.
@@ 2.4 Loss Functions during Contrastive Training
Together, these losses ensure that the model aligns LRS and XP modalities in a shared space (via contrastive loss), retains informative modality-specific content (via reconstruction), and supports spectrum translation across modalities (via prediction). In CLIP-split, this is achieved through a disentangled embedding structure with explicit shared and non-shared components, while CLIP-pr uses a unified embedding vector for all tasks. The loss curves of our CLIP-pr and CLIP-split models can be found in Appendix F.
@@ 3.2 Sample Selection for Parameter Estimation
Table 1: Comparison of Model Performance (standard deviation of the residuals $\sigma$ and coefficient of determination $R^{2}$) for different models evaluating on the held-out test datasets
@@
$\mathrm{[\alpha/Fe]}$ 0.023 / 0.872 0.021 / 0.906 0.020 / 0.912 0.020/0.913 0.020 / 0.911 0.020 / 0.916 0.020 / 0.911
$\mathrm{[C/Fe]}$ 0.041 / 0.758 0.039 / 0.792 0.037 / 0.813 0.037/0.812 0.037 / 0.812 0.037 / 0.814 0.037 / 0.813
$\mathrm{[N/Fe]}$ 0.054 / 0.598 0.052 / 0.642 0.049 / 0.664 0.049/0.665 0.049 / 0.664 0.049 / 0.667 0.049 / 0.667
$\mathrm{[Al/Fe]}$ 0.049 / 0.691 0.048 / 0.711 0.046 / 0.741 0.046/0.739 0.046 / 0.738 0.046 / 0.738 0.046 / 0.736
$\mathrm{[Ca/Fe]}$ 0.032 / 0.670 0.030 / 0.697 0.029 / 0.719 0.029/0.720 0.029 / 0.721 0.029 / 0.723 0.029 / 0.714
$\mathrm{[Mg/Fe]}$ 0.031 / 0.866 0.032 / 0.871 0.031 / 0.882 0.031/0.882 0.031 / 0.881 0.031 / 0.883 0.031 /0.880
$\mathrm{[Si/Fe]}$ 0.029 / 0.776 0.029 / 0.803 0.028 / 0.813 0.028/0.813 0.028 / 0.813 0.028 / 0.812 0.028 / 0.812
$\mathrm{[Ti/Fe]}$ 0.061 / 0.492 0.058 / 0.532 0.056 / 0.550 0.055/0.551 0.056 / 0.551 0.055 / 0.552 0.056 / 0.551
$\mathrm{[Mn/Fe]}$ 0.033 / 0.761 0.032 / 0.780 0.031 / 0.796 0.031/0.798 0.031 / 0.797 0.031 / 0.798 0.031 / 0.793
$\mathrm{[Ni/Fe]}$ 0.027 / 0.426 0.026 / 0.454 0.025 / 0.490 0.025/0.486 0.025 / 0.486 0.025 / 0.488 0.025 / 0.485
$\mathrm{[O/Fe]}$ 0.051 / 0.698 0.050 / 0.722 0.049 / 0.729 0.049/0.730 0.049 / 0.728 0.048 / 0.730 0.049 / 0.729
$\mathrm{[Cr/Fe]}$ 0.081 / 0.177 0.076 / 0.225 0.074 / 0.242 0.075/0.240 0.075 / 0.239 0.075 / 0.239 0.075 / 0.232
@@
$\Delta\nu$ 1.372/0.901 1.705/0.958 1.491/0.841 1.515/0.859 1.630/0.818 1.491/0.862 1.507/0.963
$\nu_{\rm max}$ 20.470/0.676 23.597/0.330 22.590/0.171 23.738/0.296 22.149/0.615 22.136/0.606 23.822/0.623
Mass ($M_{\odot}$) 0.095/0.518 0.094/0.570 0.086/0.674 0.087/0.670 0.084/0.669 0.085/0.664 0.089/0.658
Radius ($R_{\odot}$) 0.604/0.873 0.708/0.859 0.738/0.853 0.728/0.853 0.723/0.843 0.737/0.847 0.713/0.880
Age (Gyr) 1.565/0.655 1.488/0.684 1.397/0.721 1.347/0.744 1.347/0.751 1.352/0.747 1.337/0.723
$\Delta\Pi$ 28.078/0.891 25.703/0.904 21.471/0.923 21.814/0.920 22.057/0.917 21.346/0.924 22.713/0.925
@@
$\mathrm{[\alpha/Fe]}$ 0.103 / --0.047 0.056 / 0.737 0.049 / 0.774 0.048/0.777 0.049 / 0.773 0.049 / 0.770 0.050 / 0.765
$\mathrm{[C/Fe]}$ 0.194 / 0.073 0.127 / 0.527 0.118 / 0.547 0.118/0.553 0.118 / 0.550 0.117 / 0.549 0.118 / 0.551
$\mathrm{[N/Fe]}$ 0.115 / --4.040 0.077 / 0.643 0.072 / 0.673 0.072/0.676 0.073 / 0.672 0.072 / 0.674 0.073 / 0.669
@@
$E(BP-RP)$ 0.077 / 0.725 0.036 / 0.921 0.036 / 0.926 0.035/0.927 0.036 / 0.925 0.035 / 0.926 0.035 / 0.929
Number of wins (best $\sigma$ or $R^{2}$) 0 0 19 24 15 29 19
@@ 4.1 Model Comparison
In particular, if checked more carefully, models with in-modal reconstruction decoders (e.g., CLIP-r and CLIP-pr compared with CLIP, and CLIP-pr compared with CLIP-p) generally show improved performance, as revealed by the “number of wins” in the table (last row), fairly accounting for the MLP-based downstream models only. This indicates enhanced informativeness of the learned representations for downstream parameter estimation. Overall, these results suggest that the inclusion of in-modal reconstruction decoders improves representation quality in most downstream applications.
We hypothesize that these performance gains stem from the model’s ability to retain both shared and modality-specific (non-shared) information. In the CLIP-pr variant, this is encouraged by jointly optimizing the contrastive loss, cross-modal prediction loss, and in-modal reconstruction loss. This architecture implicitly encourages the embeddings to retain complementary information from each modality.
@@
CLIP-split achieves the highest similarity score overall, even surpassing the baseline CLIP model, possibly aided by its lower embedding dimensionality, which tends to produce higher cosine similarities. It also delivers the best performance on XP $\rightarrow$ LRS spectrum prediction, demonstrating the model’s robustness across modalities.
@@
Figure 2: Comparison between the LAMOST catalog and SpecCLIP models (including the pre-trained LRS model and the LRS branch of the CLIP-split model). From top to bottom: The radial velocity (RV) comparison as a function of the GALAH labels; [Fe/H] comparison as a function of the DESI labels; [Fe/H] comparison as a function of the GALAH labels; and [Fe/H] comparison as a function of the GALAH labels with input spectra shifted to the rest frame using the predicted RVs from the corresponding models in the top row. For RV, which is inferred using the SBI downstream model, the pre-trained LRS model and CLIP-split model have slightly larger scatter but smaller bias, compared with the LAMOST catalog; For [Fe/H], inferred using MLP downstream models (as with all other figures), the pre-trained LRS model and CLIP-split model gives either smaller scatter over the metal-poor region (referring to DESI labels) or overall smaller scatter and bias (referring to GALAH labels). The RV-corrected spectra result in similar [Fe/H] prediction performance, suggesting that the trained MLP models are relatively robust to modest Doppler shifts in the LAMOST LRS spectra. The dashed lines are the one-to-one lines. The numbers in the upper left of each panel are the mean offsets and standard deviation of the residuals (y-axis minus x-axis).
@@ 4.2.2 Gaia XP
Figure 5: Two examples of in-modal retrieval, cross-modal retrieval, cross-modal prediction, and the LAMOST (Gaia) spectra corresponding to Gaia (LAMOST) in-modal retrieval. The similarity scores are defined in the projected embedding space.
Table 3: Comparison of Model Performance (standard deviation of the residuals $\sigma$ and coefficient of determination $R^{2}$) for pre-trained XP models with different embedding dimensions (256, 343, 512, 768), where 768 is adopted in this paper
@@ 4.3.1 Spectral Retrieval
Figure 5 shows examples of in-modal and cross-modal retrievals given a specific query spectrum. Retrieval is based on cosine similarity in the (projected) embedding space, using either the full or shared embeddings (for CLIP-split). In this figure, the search is performed using the CLIP-pr model on a test set of 82,057 spectra, with the query spectrum excluded from the database. In both LAMOST LRS and Gaia XP cases, the retrieved spectra closely resemble the query spectra, indicating that the model has learned well-aligned representations across modalities.
In practice, additional strategies can be used to retrieve spectra using auxiliary catalog links. For example, given a query LAMOST spectrum and a LAMOST-to-Gaia cross-match library, one could first retrieve the top LAMOST matches (in-modal) and then fetch their Gaia counterparts via the library. Alternatively, if the paired Gaia spectrum of the query is known, one could retrieve similar Gaia spectra directly (in-modal), or indirectly by performing a Gaia-to-LAMOST retrieval followed by a database lookup to obtain the corresponding Gaia spectra. While Figure 5 presents only two retrieval use cases, assuming that we know only the information of the query spectrum itself, not its paired other-modal spectrum. The other more elaborate approaches are straightforward extensions.
@@ 4.3.2 Spectral Prediction
SpecCLIP’s cross-modal decoders also support spectral translation, that is, predicting the spectrum in one modality from a spectrum in another. These decoders operate directly on the projected embeddings (shared embeddings in the case of CLIP-split), using learned mappings between modalities.
@@
However, for a subset of sources—particularly in the LRS $\rightarrow$ XP direction—prediction quality deteriorates. Apart from the potential effect of extinction poorly-learned by the trained model, this discrepancy may suggest that the source does not follow the behavior of a typical single star. For instance, it could be an unresolved binary or an otherwise anomalous object. These cases highlight an exciting future direction, using cross-modal prediction error as a basis for anomaly detection.
@@ 5 Discussion
5.1 SBI Performance for [Fe/H] and Radial Velocity
In our experiments with SBI, we observed that applying NSF to projected embeddings led to degraded performance in the low-iron abundance regime ($\mathrm{[Fe/H]}\lesssim-2$). Similarly, radial velocity predictions from LAMOST LRS exhibited systematic underestimation of the absolute value at large radial velocities; a similar issue broadly discussed in machine learning-based estimators (Ting, 2024).
@@ 5.1 SBI Performance for [Fe/H] and Radial Velocity
A common assumption is that compression improves downstream performance. However, our Gaia XP foundation model shows a more nuanced behavior. In Table 3, we vary the embedding dimension and find that reducing it from 343 (the original XP spectrum length) to 256 leads to comparable performance. In contrast, using embedding dimensions equal to or greater than the input size yields better results.
@@ 5.5 Training Sample Size and Dataset Configuration
Nevertheless, even with this moderate sample size, our models match or outperform state-of-the-art results in the literature (Section 4.2). To maximize data usage, we report metrics based on held-out test sets, while for plots generated using the downstream models, we combine the training and validation sets for training, with early stopping based on performance on the test set.
@@ 5.7 Fine-tuning Loss Weights
Table 4: Comparison of Model Performance (standard deviation of the residuals $\sigma$ and coefficient of determination $R^{2}$) between masked transformer (MT) and MLP-based ordinary auto-encoder (OAE)
@@
$\mathrm{[\alpha/Fe]}$ 0.023 / 0.872 0.021 / 0.906 0.020 / 0.904
$\mathrm{[C/Fe]}$ 0.041 / 0.758 0.039 / 0.792 0.039 / 0.776
$\mathrm{[N/Fe]}$ 0.054 / 0.598 0.052 / 0.642 0.053 / 0.624
$\mathrm{[Al/Fe]}$ 0.049 / 0.691 0.048 / 0.711 0.049 / 0.693
$\mathrm{[Ca/Fe]}$ 0.032 / 0.670 0.030 / 0.697 0.031 / 0.688
$\mathrm{[Mg/Fe]}$ 0.031 / 0.866 0.032 / 0.871 0.030 / 0.873
$\mathrm{[Si/Fe]}$ 0.029 / 0.776 0.029 / 0.803 0.029 / 0.793
$\mathrm{[Ti/Fe]}$ 0.061 / 0.492 0.058 / 0.532 0.059 / 0.507
$\mathrm{[Mn/Fe]}$ 0.033 / 0.761 0.032 / 0.780 0.033 / 0.758
$\mathrm{[Ni/Fe]}$ 0.027 / 0.426 0.026 / 0.454 0.027 / 0.445
$\mathrm{[O/Fe]}$ 0.051 / 0.698 0.050 / 0.722 0.051 / 0.704
$\mathrm{[Cr/Fe]}$ 0.081 / 0.177 0.076 / 0.225 0.079 / 0.200